Cutting a yard must be done in three steps, cutting, trimming, and blowing. If one person does this, it takes 90 minutes to do one yard. If two people divide the work, the first person takes 60 minutes to cut. The seconnd person takes 45 minutes to trim and blow. (These jobs can be performed simultaneously). A) How many yards can the single person do in an eight-hour day? B) How many yards can the team of two do in an eight-hour day?
Answer #1

A)

One person takes 90 minutes for 1 yard.

Total Yard can be done in 8 hours = (8*60)/90 = 5.33 yards

B)

Time to cut one yard = 60 min

Time to trim and blow = 45 minutes

Since processing time of yard in higher hence it will be bottleneck.

Capacity of Team = (8*60)/60 = 8 yards